OpenStudy (anonymous):

solve for x x+2/x = 3 solve for x x+2/x = 3 @Mathematics

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

from ur equation, u get x+2=3x, then u get 2x=2, so x=1

OpenStudy (turingtest):

\[x+{2\over x} = 3\] or\[ {x+2\over x}=3\]?????????

OpenStudy (turingtest):

(umichzane solve the second equation above)

OpenStudy (anonymous):

its the first one.. and there should like be two solutions

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Oops, ur right... x=1 or x=2, it becomes a quadratic that can be factored and solved

OpenStudy (turingtest):

multiply it all be x first and you get x^2+2=3x x^2-3x+2=0 (x-2)(x-1)=0 x=(2,1)

OpenStudy (turingtest):

they usually like to write it as x=(1,2) but whatever (lowest root to highest)

OpenStudy (anonymous):

ahh thank you so so much!! that makes sence!!! :D

OpenStudy (turingtest):

they usually like to write it as x=(1,2) (lowest to highest root), but whatever